OpenStudy (anonymous):

Solve x2 – 8x + 4 = 0

OpenStudy (anonymous):

x^2-8x+4=0 x^2-8x=-4 x^2-8x+16= -4+16 (x-4)^2=12 x-4=±√(12) x=4±√(4*3) x=4±2√(3) x=4+2√(3), 4-2√(3) x=7.46, 0.54

OpenStudy (anonymous):

if u knew the answer, why'd u ask this?

OpenStudy (anonymous):

I was typing the possible answers..........................................

OpenStudy (anonymous):

ah, i didnt know it was multiple choice